OpenStudy (anonymous):

need help with this math problem. Describe and correct the error in the solution. 4d+9d-(7-6d)=4d+9d-7-6d =4d+9d-6d-7 =7d-7 need help with this math problem. Describe and correct the error in the solution. 4d+9d-(7-6d)=4d+9d-7-6d =4d+9d-6d-7 =7d-7 @Mathematics

OpenStudy (anonymous):

YOu forgot to distibute the negative sign

OpenStudy (anonymous):

4d+9d+7+6d

OpenStudy (anonymous):

19d+7

OpenStudy (anonymous):

could you possibly walk my daughter through how you got that,, im not to good at math and shes not understanding what you mean? sorry its confusing to us. lol

OpenStudy (anonymous):

4d+9d-(7-6d) = 4d+9d-7+6d = 4d+9d+6d-7 = 19d - 7

OpenStudy (anonymous):

Do you see the negative sign before the bracket? To make it simpler put a -1 in front:4d+9d-1(7+6d) Now you gotta do distribution 4d+9d-1*7-1*-6d Now simplify: 4d+9d-7+6d 19d-7

OpenStudy (anonymous):

Is that clear?

OpenStudy (anonymous):

yes thank you very much